OpenStudy (anonymous):

A desk top has the shape of a trapezoid. The two parallel sides have lengths 2 m and 2.8m, and the distance between the sides is 1.5m. To protect the wood surface, a blotter will be made, at a cost of$25 per square meter, to cover the entire desk top. Find the total cost of the blotter.

OpenStudy (anonymous):

a=altitude, b=base, a trapezoid has 2 bases. a((b[1]+b[2])/2)=A =1.5((2+2.8)/2) =1.5(4.8/2) =1.5*2.4 =3.6 m^2 3.6*25=$90. Cost of the blotter .

OpenStudy (anonymous):

metal please